# BAGELS: Benchmarking the Automated Generation and Extraction of Limitations from Scholarly Text

### v1

@@ Abstract
In scientific research, “limitations” refer to the shortcomings, constraints, or weaknesses within a study. A transparent reporting of such limitations can enhance the quality and reproducibility of research and improve public trust in science. However, authors often a) underreport them in the papers’ text and b) use hedging strategies to satisfy editorial requirements at the cost of readers’ clarity and confidence. This underreporting behavior, along with an explosion in the number of publications, has created a pressing need to automatically extract/generate such limitations from scholarly papers. In that direction, we report a complete architecture for computational analysis of research limitations. Specifically, we a) create a dataset of limitations in ACL, NeurIPS, and PeerJ papers by extracting them from papers’ text and integrating them with external reviews; b) propose methods to automatically generate them using a novel Retrieval Augmented Generation (RAG) technique; c) create a fine-grained evaluation framework for generated limitations and provide a meta-evaluation for the proposed evaluation techniques. Code and datasets are here: github | huggingface
[…]
@@
Ibrahim Al Azher, Miftahul Jannat Mokarrama, Zhishuai Guo, Sagnik Ray Choudhury, Hamed Alhoori Northern Illinois University, Dekalb, IL, USA University of North Texas, Denton, TX, USA {iazher, mmokarrama1, zguo, alhoori}@niu.edu, sagnik.raychoudhury@unt.edu
@@ 1 Introduction
Figure 1: System Architecture. We build a dataset by extracting limitations from a paper’s text (when mentioned) and reviews, build a generator for papers that do not explicitly mention limitations, and develop a novel framework for reference-based evaluation of generated limitations.
@@
In the context of a scientific article, “limitations” refer to the inherent shortcomings, constraints, or weaknesses within a study that may influence the results or limit the generalizability of its findings Ross and Bibler Zaidi (2019). These limitations can arise from various elements of the research process, including the methodology, theoretical framework, data collection, experimentation, and analysis Ioannidis (2007). Limitations explicitly acknowledged by the authors often include internal validity issues, measurement errors, potential confounding factors, and the failure to measure important variables Puhan et al. (2009).
@@
Openly discussing limitations is crucial for several reasons. It ensures the credibility and ethical standards of a scientific inquiry Bunniss and Kelly (2010); Chasan-Taber (2014); Annesley (2010). By clearly defining the boundaries of a study, researchers provide a realistic context for their findings, supporting accurate interpretation, transferability, and reproducibility Ioannidis (2007); Eva and Lingard (2008). Acknowledging limitations fosters scientific integrity by demonstrating a commitment to honesty Žydžiūnaitė (2018). It also opens pathways for refining research methods, opening areas for future work Azher et al. (2025), and leading to innovations that help future researchers avoid repeating the same shortcomings Escande et al. (2016).
@@
Despite the benefit of honest admission of such limitations, researchers sometimes remain reluctant to include them in the text or articulate them in detail Ioannidis (2007); Ter Riet et al. (2013). Concerns about the potential impact on publication chances and, subsequently, career progression Montori et al. (2004) can encourage this biased behavior. When forced to explicitly acknowledge limitations in the text [^1] As is the practice in the NLP and ML community in recent years., authors sometimes resort to using generic or irrelevant terms, which do not meaningfully inform the reader about the study’s real constraints Ross and Bibler Zaidi (2019). Moreover, limitations can sometimes serve as a form of “hedging”, where authors present their findings cautiously to avoid making definitive statements Hyland (1998). This approach, while safer for the authors, does not contribute to the clarity or usefulness of the research.
@@
Failure to disclose limitations undermines the scientific process and misleads readers, reviewers, and policymakers, potentially preventing them from recognizing the constrained nature of findings and possible biases Greener (2018). NLP research in automatic extraction, generation, and evaluation of limitations from scholarly papers is limited due to the lack of standardized datasets, novel methods, and evaluation framework. This study takes a step toward bridging this gap.
The number of scientific publications has exploded recently Larsen and Von Ins (2010). That, combined with the problems of self-reporting, has necessitated a need for computational methods to study the limitations of scholarly work. In that direction, our contributions are as follows (see Figure 1):
  • We create a dataset for limitations reported in research articles. For each paper, we extract the limitations from the papers’ text and OpenReview feedback  [^2] https://openreview.net/. This creates a valuable benchmark for future studies that want to extract and analyze limitations from research papers and study their impact on scholarly progress. Notably, due to the integration of authored text and external data, this benchmark suffers less from the self-reporting bias, providing a broader and more nuanced perspective about limitations.
  • We create a novel RAG system to generate limitations. This is a useful tool for editors or researchers in the under-reporting or hedging scenario.
  • Evaluating the generated limitations is a key challenge. Many traditional metrics for measuring text generation quality depend on lexical overlap (ROUGE Lin (2004), BLEU Papineni (2001)) or semantic similarity (BERTScore Zhang et al. (2019), MoverScore Zhao et al. (2019)). They are not sufficient for our purpose as they overprioritize common terms (e.g., tokens such as “bias”, “dataset”, and “generalizability” are standard in the limitations section of most AI/ML/NLP papers and should not affect the evaluation score). We develop a new evaluation framework that uses LLMs-as-judges to provide fine-grained comparisons and detailed error analysis, thus offering a more transparent and actionable evaluation for limitations.
@@ 2 Related Work
Various studies have studied the reporting, or lack thereof, of ‘limitations’ in scholarly papers. Ioannidis (2007) found that only 17% of top-tier journal articles mentioned limitations, with just 1% in abstracts. Similarly, Puhan et al. Puhan et al. (2012) identified 27% of biomedical papers lacked limitations, risking overestimated research reliability. Moreover, Goodman Goodman et al. (1994) found that in the peer review process acknowledging limitations is often problematic. Few journals require discussing limitations Ioannidis (2007), risking biased reviews and weakening scientific dialogue Horton (2002), necessitating reforms for transparency.
These gaps underscore the need for a systematic framework for limitations generation, an area that remains underexplored. Faizullah et al. Faizullah et al. (2024) propose an LLM-chain pipeline to summarize and iteratively refine candidate limitations; Al Azher et al. (2024) integrates topic modeling with LLMs to derive and elaborate structured limitation themes; and Azher (2024) develops a graph-augmented LLM approach for generating comprehensive limitation statements. Other work focuses on overcoming limitations of charts and graph by generating meanigful captions Al Azher and Alhoori (2024). However, these studies are confined to ACL/EMNLP corpora and author-stated limitations, overlooking reviewer comments (e.g., OpenReview), and evaluate only with ROUGE, BERTScore, and coherence metrics, which do not assess finer-grained topic or contextual alignment. Our proposed framework addresses existing limitations by incorporating NeurIPS papers alongside ACL, collecting OpenReview comments for each NeurIPS paper to establish a broader ground truth, and implementing a limitation-level text evaluation method to preserve granularity.
@@
Evaluating NLP outputs is crucial for assessing quality, accuracy, and relevance. Traditional metrics like ROUGE and BLEU struggle with semantics, while BERTScore improves similarity but relies on references and lacks meaningful error analysis. Recent advancements in large language models (LLMs) have revolutionized text evaluation methods Zheng et al. (2023) using zero-shot and in-context learning Wei et al. (2022). For example, GPTScore Fu et al. (2023) uses LLMs to evaluate how likely a text could be generated based on a given context, TIGERScore Jiang et al. (2023) provide explainable error analysis, and PandaLM Wang et al. (2023) distinguish superior models among multiple models. Other approaches, such as AttrScore Yue et al. (2023), evaluate whether a reference supports or contradicts generated statements. SummacConv Laban et al. (2022) breaks documents into sentence-level units and removes summary sentences with low entailment probability.
Despite their promise, LLM-based evaluations have issues such as positioning bias, where altering the order of inputs affects results. To mitigate positioning bias, we systematically alter the position of limitations and select outputs that remain consistent across different orderings. Moreover, while these text evaluation methods focus on lexical overlap and surface-level LLM-based analysis, our proposed approach advances beyond these limitations by uniquely combining granularity-aware evaluation, considering LLM as a judge, quantitative scoring, with topic-level agreement.
3 Limitation Extraction
Our limitation extraction is twofold. At first, we extracted limitations with a science parse  [^3] https://github.com/allenai/science-parse tool. If a paper has a separate limitation section, we were able to extract it using this tool. But if the limitation has no separate section, we used Python regex to check the word ‘limitation’ and extracted all consecutive texts before another section appeared. However, this approach also collected some noisy sentences unrelated to the limitation. So, in the later part, we applied LLM and sent the texts to extract only limitations. We also used Selenium to collect reviews from OpenReview. Later, the ‘author mentioned limitation’ is sent to the zero-shot LLM to extract limitations, and the review from ’OpenReview’ is sent to the zero-shot LLM to extract shortcomings/weaknesses. Through this process, we removed the noisy sentences. (§ 3.1)
@@ 3.1 Dataset of Extracted Limitations
Granularity. A key challenge in constructing a dataset of research limitations is defining the appropriate level of granularity. Should a limitation be captured as a single phrase, a full sentence, or an entire paragraph? We define a limitation as a sequence of sentences, as individual sentences often do not encapsulate multiple limitations. In contrast, a single limitation can extend across multiple sentences, sometimes forming a complete paragraph.
@@
Extraction Sources. Two primary sources form the basis of our dataset: (1) limitations explicitly acknowledged by authors, and (2) those highlighted through peer-review commentary. Although author-reported limitations often provide well-structured insights, previous research indicates that such limitations may be underreported or carefully hedged. To address this gap, we incorporate OpenReview [^4] https://openreview.net/ comments, in which peer reviewers frequently identify additional constraints or weaknesses not mentioned by the authors.
@@
Our dataset consists of papers from conferences in natural language processing and machine learning: ACL [^5] https://aclrollingreview.org/cfp, NeurIPS [^6] https://neurips.cc/public/guides/PaperChecklist, and PeerJ  [^7] https://peerj.com/benefits/indexing-and-impact-factor/. We include 6,932 NeurIPS papers from 2021 and 2022 and 5,739 ACL papers from 2023-2024, 1000 biological and medical science papers from PeerJ. In addition, we integrate OpenReview comments for 2,802 papers from NeurIPS, thus capturing both self-reported limitations and those externally identified by reviewers. All of the PeerJ paper contains self-reported Limitations alongside other sections and peer review comments. The first experiment evaluates state-of-the-art LLMs’ ability to extract limitations from papers (when they are explicitly mentioned) and their reviews. For each paper, we use GPT-4o mini to extract and get an average of $8$ limitations from a paper and $10$ from open review.
LLM as a Ground Truth Limitation Extractor. Identifying and isolating limitations can be challenging, as they often appear in sections such as Conclusion, Discussion, or Future Work and may also be dispersed throughout OpenReview feedback. To address this, we implement a targeted extraction process designed to be both efficient and accurate. Rather than processing entire papers, which would be computationally expensive due to larger context windows, we first apply Python regex to identify and extract relevant sections. Some research papers include a dedicated limitations section; we extract the text of this section using the AllenAI Science Parse tool [^8] https://github.com/allenai/science-parse. When a paper doesn’t have such a well-defined section, we identify sentences containing the word limitations and continue extracting subsequent sentences until we encounter one of the following sections: Conclusion, Future Work, Ethics, Acknowledgement, or Grant. Then, we extracted the limitations from the research paper by LLM to remove noisy sentences and make segments.
To capture broader perspectives from peer reviews, we aggregate comments from multiple OpenReview responses into a single text. Finally, we employ GPT-4o  [^9] https://openai.com/index/hello-gpt-4o/ (Figure 4) to segment the extracted text and identify distinct limitation statements, distinguishing those reported by authors from those highlighted by reviewers. We evaluate the quality of these extracted limitations through a user study (§5).

### v2

@@ Abstract
In scientific research, “limitations” refer to the shortcomings, constraints, or weaknesses of a study. A transparent reporting of such limitations can enhance the quality and reproducibility of research and improve public trust in science. However, authors often underreport limitations in their papers and rely on hedging strategies to meet editorial requirements at the expense of readers’ clarity and confidence. This tendency, combined with the surge in scientific publications, has created a pressing need for automated approaches to extract and generate limitations from scholarly papers. To address this need, we present a full architecture for computational analysis of research limitations. Specifically, we (1) create a dataset of limitations from ACL, NeurIPS, and PeerJ papers by extracting them from the text and supplementing them with external reviews; (2) we propose methods to automatically generate limitations using a novel Retrieval Augmented Generation (RAG) technique; (3) we design a fine-grained evaluation framework for generated limitations, along with a meta-evaluation of these techniques. Code and datasets are available at: Code: https://github.com/IbrahimAlAzhar/BAGELS_Limitation_Gen Dataset: https://huggingface.co/datasets/IbrahimAlAzhar/limitation-generation-dataset-bagels
Accepted to the Findings of the 2025 Conference on Empirical Methods in Natural Language Processing (EMNLP 2025).
@@
Ibrahim Al Azher, Miftahul Jannat Mokarrama, Zhishuai Guo, Sagnik Ray Choudhury, Hamed Alhoori Northern Illinois University, DeKalb, IL, USA University of North Texas, Denton, TX, USA {iazher1, mmokarrama1, zguo, alhoori}@niu.edu, sagnik.raychoudhury@unt.edu
@@ 1 Introduction
Figure 1: System architecture for dataset creation, limitation generation, and evaluation.
@@
In scientific articles, “limitations” refer to the inherent shortcomings, constraints, or weaknesses of a study that may influence its results or restrict the generalizability of its findings Ross and Bibler Zaidi (2019). Such limitations can arise from various aspects of the research process, including the methodology, theoretical framework, data collection, experimentation, and analysis Ioannidis (2007). Authors commonly acknowledge issues such as internal validity concerns, measurement errors, confounding factors, and the omission of important variables Puhan et al. (2009).
@@
Openly discussing limitations is crucial. It upholds credibility and scientific integrity by demonstrating a commitment to ethical and transparent research practices Bunniss and Kelly (2010); Chasan-Taber (2014); Annesley (2010); Žydžiūnaitė (2018). It also clarifies the scope of a study, supporting accurate interpretation, transferability, and reproducibility Ioannidis (2007); Eva and Lingard (2008). In addition, it helps researchers avoid repeating the same shortcomings Escande et al. (2016) while creating opportunities to refine methods and guide future research Azher et al. (2025).
@@
Despite these benefits, researchers are often reluctant to include limitations or articulate them in detail Ioannidis (2007); Ter Riet et al. (2013). Concerns about the potential impact on publication chances and career progression Montori et al. (2004) can reinforce this tendency. Even when required to acknowledge limitations, as is now common in NLP/ML research, authors sometimes resort to generic or irrelevant statements that obscure the study’s real constraints Ross and Bibler Zaidi (2019). Moreover, limitations may serve as a form of hedging, where findings are presented cautiously to avoid making definitive claims Hyland (1998). This practice, while safer for authors, reduces the clarity and usefulness of the research.
@@
Failure to disclose limitations undermines the scientific process and misleads readers, reviewers, and policymakers, preventing recognition of constrained findings and potential biases Greener (2018). Meanwhile, the volume of scientific publications has surged Bornmann et al. (2021). These factors highlight the need for computational methods to study research limitations. However, progress in NLP toward automatic extraction, generation, and evaluation of limitations remains limited, largely due to the lack of standardized datasets, novel methods, and robust evaluation frameworks. This study takes a step toward closing this gap.
Our contributions are as follows (see Figure 1):
  • Dataset creation. We build a dataset of research limitations by extracting them from papers and their reviews. By integrating author-reported and reviewer-identified limitations, this benchmark reduces self-reporting bias and provides a broader, more reliable resource for analyzing limitations and their impact on research.
  • Limitation generation. We design a novel RAG system to automatically generate limitations, offering a way to supplement papers with high-quality, context-aware limitation statements.
  • Evaluation framework. We introduce a new evaluation paradigm for generated limitations. Unlike traditional metrics (e.g., ROUGE Lin (2004), BLEU Papineni (2001), BERTScore Zhang et al. (2019), MoverScore Zhao et al. (2019)), which overemphasize common terms (e.g., bias, dataset, and generalizability), our framework leverages LLMs-as-judges for fine-grained, interpretable assessments and actionable error analysis.
@@ 2 Related Work
Several studies have examined how limitations are reported in papers. Ioannidis (2007) found that only 17% of top-tier articles mentioned limitations, with just 1% doing so in abstracts. Similarly, Puhan et al. (2012) reported that 27% of biomedical papers lacked limitations, risking overestimation of research reliability. Goodman et al. (1994) noted that acknowledging limitations is often problematic in peer review. Few journals require discussing limitations Ioannidis (2007), which can bias reviews and weaken scientific dialogue Horton (2002), highlighting the need for greater transparency.
Recent work has explored computational approaches to research limitations. Faizullah et al. (2024) proposed an LLM-chain pipeline to summarize and refine candidate limitations. Al Azher et al. (2024) integrated topic modeling with LLMs to derive structured limitation themes. Al Azher (2024) developed a graph-augmented LLM method for generating detailed limitation statements. Other studies address the shortcomings of visualizations by generating more meaningful captions for charts and graphs Al Azher and Alhoori (2024). However, these studies are limited to ACL/EMNLP corpora and rely on author‐stated limitations, and use metrics such as ROUGE and BERTScore that miss finer-grained contextual alignment. Concurrent to our work, Xu et al. (2025) also address these gaps by introducing LIMITGEN, a benchmark that incorporates human-written peer reviews to systematically evaluate how well LLMs identify limitations. Our framework complements this effort by additionally leveraging cited papers for broader context and introducing a novel limitation-level evaluation method to preserve granularity.
@@
Evaluating NLP outputs is essential for assessing quality, accuracy, and relevance. Traditional metrics like ROUGE and BLEU struggle with semantics, while BERTScore improves similarity but relies on references and lacks meaningful error analysis. Advances in large language models (LLMs) have opened new evaluation avenues Zheng et al. (2023), from zero-shot and in-context learning Wei et al. (2022) to specialized approaches such as GPTScore Fu et al. (2023), TIGERScore Jiang et al. (2023), and PandaLM Wang et al. (2023). Other methods include AttrScore Yue et al. (2023), which checks factual support, and SummacConv Laban et al. (2022), which filters low-entailment sentences. Despite their promise, LLM-based evaluations face issues such as positioning bias, where input order can shift results. We address this by randomizing order and retaining stable outputs. More broadly, our evaluation advances beyond prior work by combining granularity-aware scoring, topic-level agreement, and LLMs-as-judges.
Taken together, prior research shows both the need and the opportunity for a more systematic treatment of research limitations. Building on these insights, our work unifies dataset construction, limitation generation, and evaluation into a single framework, laying the foundation for more transparent and reproducible analysis of limitations.
3 Limitation Extraction & Evaluation
@@ 3.1 Dataset of Extracted Limitations
Granularity. A key challenge in building a dataset of research limitations is defining the appropriate level of granularity. Should a limitation be captured as a single phrase, a full sentence, or an entire paragraph? We define a limitation as a sequence of sentences, as individual sentences often do not encapsulate multiple limitations. In contrast, a single limitation can extend across multiple sentences, sometimes forming a complete paragraph.
@@
Extraction Sources. Two primary sources form the basis of our dataset: (1) limitations explicitly acknowledged by authors, and (2) those highlighted through peer-review commentary. Although author-reported limitations often provide well-structured insights, previous research indicates that such limitations may be underreported or carefully hedged. To address this gap, we incorporate review comments, where reviewers often highlight additional constraints or weaknesses not mentioned by the authors.
@@
Our dataset includes papers from major NLP and ML conferences, including ACL [^1] https://aclrollingreview.org/cfp and NeurIPS [^2] https://neurips.cc/public/guides/PaperChecklist, as well as biomedical research from PeerJ  [^3] https://peerj.com/benefits/indexing-and-impact-factor/. We collect 6,932 NeurIPS papers (2021-2022), 5,739 ACL papers (2023-2024), and 1000 papers from PeerJ. In addition, we integrate OpenReview  [^4] https://openreview.net/ comments for 2,802 papers from NeurIPS. All of the PeerJ papers contain self-reported limitations alongside other sections and peer review comments. For each paper, we use LLM to extract and get an average of $8$ limitations from a paper and $10$ from their reviews.
3.2 Extraction Process
We extract spans (blocks) of text from papers or review comments, and then refine them with LLMs, as opposed to passing in the entire paper to an LLM. This strikes a balance between accuracy and LLM usage cost.
1. Limitation Span Extraction: This step extracts blocks of text from the papers that correspond to limitations. We consider both explicit and implicit limitation statements:
a. Explicit limitations. These appear in a dedicated limitations section or subsection. We identify them using the AllenAI Science Parse tool  [^5] https://github.com/allenai/science-parse, which segments papers into a structured JSON format, allowing for direct and reliable extraction of these dedicated sections. For peer review content in NeurIPS papers, we used Selenium to scrape the main review field from OpenReview, which typically includes both strengths and weaknesses of a paper.
b. Implicit limitations. These are embedded in broader sections such as discussion or conclusion. To identify them, we apply a Python regex script that searches for keywords such as limitation(s), or shortcoming(s). To improve precision, we exclude sections where limitations are rarely discussed (e.g., abstract, introduction, related work). Our script begins extraction when a limitation-related keyword is detected and continues until a terminal section marker is reached; extraction stops at terms such as acknowledgements, grant, future work, discussion, conclusion, or appendix. Although this process is effective, the regex approach for implicit limitations can occasionally capture irrelevant sentences, introducing noise into the results.
2. Refinement via LLM: To improve precision, we use an LLM to filter meaningful limitations from the tool-extracted ones (from both papers and review) by removing noisy sentences. Importantly, we strictly instruct the LLMs to extract limitation statements without paraphrasing, altering, or generating new content, and producing them as a structured sequence of sentences, denoted by $L_{i}=\{l_{i1},l_{i2},\dots,l_{ix}\}$. To incorporate broader perspectives from peer reviews, we first aggregate comments from multiple review responses into a single consolidated text. We then prompt the LLMs (Figure 3, appendix) to segment this text and identify distinct limitation statements by reviewers, with the latter being denoted as $R_{i}=\{r_{i1},r_{i2},\dots,r_{ix}\}$.
Following this extraction, a master LLM is tasked with merging the author-reported limitations $L_{i}$ and the reviewer-identified limitations $R_{i}$ of input paper $P_{i}$. The model is explicitly instructed to merge only those limitation statements that were identical or semantically equivalent across both the author-mentioned limitations and the peer review. As before, the model is restricted from changing, rephrasing, or reordering any sentences during the merge process, and we get final Ground truth extracted limitations $G_{i}=\{g_{i1},g_{i2},\dots,g_{ix}\}$. Finally, we evaluate the quality of these extracted and merged limitations through a user study described in § 3.3. We use GPT 4o-mini as both the extractor and master LLM (Examples of limitations extraction by LLM from NeurIPS, ACL, and OpenReview are provided in the Appendix in Figure 6, 7, and 8, respectively).
3.3 Limitation Extraction Evaluation
Are the limitations extracted or generated? The first goal in the evaluation process is to check if the LLM extracted limitations are grounded in the text, i.e., they only come from the input (papers/reviews) and not from the LLMs’ parametric knowledge or hallucinations. For this, we employ three annotators (separate from this paper’s authors) [^6] CS graduate students with research experience in NLP and AI.
The first ground truth consists of only author-mentioned limitations. We choose a sample of 100 limitations from ACL, NeurIPS, and PeerJ, and for each, we show them the source and ask a Yes/No question, whether they thought the LLM extracted the limitation from the source without generating text. Each annotator answer positively in $90\%$ of cases (avg $\pm$ std= $95\pm 2.45\%$) (Table 1).
Model Role Sample U1 U2 U3
GPT 4o-mini Extractor 100 92 95 98
Table 1: Evaluating LLM as an extractor role with human annotator (U).
In the second evaluation, two annotators manually verified the extracted limitations from 1000 papers from NeurIPS and PeerJ and their reviews. The annotators assessed whether 1) each LLM-extracted author mentioned limitation was grounded in the source paper, (2) each extracted limitation from the peer review was also grounded in the review, and (3) the merged set (limitation + review) included only truly overlapping or matching limitations between the two sources. Their analysis confirmed that all extracted limitations were faithfully sourced, with no instances of hallucinated, noisy, or newly generated content. We also computed the performance of the Llama3 70B for this extraction task, and the result was unsatisfactory.
The quality of the extraction. The SMEs from the last step annotated 500 ACL papers and 100 NeurIPS papers: one annotator extracted limitations (taking the full section when explicit, or selecting limitation-related sentences when implicit), and two others verified the results. We then compared the tool-based (GPT-4o mini) extractions against this gold standard. Notably, the human-extracted (gold) limitations were not segmented; therefore, we combined the LLM-extracted limitations and compared them with the gold ones using cosine similarity, precision, recall, F1, and fuzzy matching [^7] These strings are tokenized. (Table 2): ACL achieved a strong F1 of 85.69, likely aided by more frequent explicit limitation sections. NeurIPS yielded a moderate F1 of 72.42, reflecting the more scattered, implicit presentation of limitations where LLM should be utilized to remove noisy information.
Dataset CS P R F1 Fuzzy
ACL 89.38 89.63 84.93 85.69 91.18
Neurips 78.08 68.76 84.13 72.42 70.26
Table 2: Performance between Human Extracted Limitations vs Tool Extracted Limitations in Cosine Similarity (CS), Precision (P), Recall (R), F1 score (F1), and Fuzzy matching
3.4 Dataset Applications